# D-Judge: How Far Are We? Accessing the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

### v2

@@ Appendix A Appendix Overview
Appendix B Details of the D-ANI Dataset
B.1 Overview
The D-ANI dataset is a carefully curated collection of both natural and AI-generated images, designed to support comprehensive evaluations of image generation models. The natural images are sourced from the COCO validation set. AI-generated images were produced using eight different models. These images were generated under three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). Representative samples from the D-ANI dataset are shown in Figures 9,  10 and  11. For a detailed breakdown of the D-ANI dataset composition, please refer to Table 6.
Figure 9: Examples of AI-generated content (AIGC) created using Text-to-Image (T2I) guidance.
Figure 10: Examples of AI-generated content (AIGC) created using Image-to-Image (I2I) guidance.
Figure 11: Examples of AI-generated content (AIGC) created using Text-and-Image-to-Image (TI2I) guidance.
B.2 Natural Images
The D-ANI dataset features a carefully curated selection of natural images sourced from the COCO validation set. This subset comprises 5,000 images, each paired with at least five unique captions, totaling 25,014 captions. These rich and varied textual descriptions enhance the dataset’s utility across a wide range of image-to-text and text-to-image tasks, making it an invaluable resource for evaluating and training image generation and understanding models. In this study, we use these images and their accompanying (five) captions to guide the generative models in creating new images.
B.3 Generated Images
The D-ANI dataset includes AI-generated images created using several advanced generative models across three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). These modes collectively contributed a significant volume of images to the dataset. The models utilized in this process include DF-GAN, GALIP, various versions of Stable Diffusion (v1.4, v1.5, v2.1, XL), Versatile Diffusion, DALL·E 2 and DALL·E 3. Each model, except DALL·E 2 and DALL·E 3, generated 25,000 images per guidance mode. DALL·E 2 and DALL·E 3 produced 5,000 images for per mode, resulting in a total of 440,000 generated images. It is important to note that DF-GAN, GALIP and DALL·E 3 were used only for T2I, while DALL·E 2 was employed for both T2I and I2I modes. For detailed information, please refer to Table 6.
@@ B.3 Generated Images
  • Versatile Diffusion: (Xu et al., 2023b) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution. Its versatility lies in its ability to adapt the diffusion process to the specific requirements of each task, making it suitable for diverse applications.
@@
B.4 Image Resolution
The resolution of the generated images varies across models. GALIP produces images at 224x224 pixels, while DF-GAN generates slightly higher-resolution images at 256x256 pixels. Models such as various versions of Stable Diffusion and DALL·E 2 generate images at a resolution of 512x512 pixels. Notably, DALL·E 3 delivers the highest resolution among these models, producing images at 1024x1024 pixels.
B.5 Composition
The D-ANI dataset offers a robust foundation for evaluating the performance and capabilities of various AIGC models under different conditions and guidance modes. By encompassing a broad spectrum of images and resolutions, the dataset ensures comprehensive coverage of the potential use cases and challenges associated with AI-generated content. The combination of natural and generated images, alongside diverse textual descriptions, facilitates a thorough assessment of image generation models, supporting the development of more accurate and reliable AI systems.
Table 6: The details of the D-ANI dataset.
Types DFGAN GALIP SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 25,000 25,000 25,000 25,000 25,000 25,000 25,000 5,000 5,000
I2I - - 25,000 25,000 25,000 25,000 25,000 5,000 -
TI2I - - 25,000 25,000 25,000 25,000 25,000 - -
@@ C.1 Overview
To evaluate image quality, we divided the assessment into three sub-aspects: pixel-level similarity, frame-level visual metrics, and visual content distribution.
  • Pixel-level: Metrics such as SSIM, PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity. These metrics measure how closely the generated images match natural ones in terms of structural integrity and perceptual similarity.
@@
For alignment, we utilize the CLIP Score (Radford et al., 2021a) to measure the correspondence between text descriptions and the generated images. This metric evaluates how well the generated content aligns with the given prompts by embedding both text and image into a shared space and calculating their similarity.
@@
C.2 Naive Image Quality
We evaluated the naive image quality of AI-generated images versus natural images using a range of metrics. These metrics assess both pixel-level and frame-level qualities, as well as visual content distribution, to capture the holistic differences between AI-generated and natural images.
Pixel-Level Quality Metrics:
@@ C.2 Naive Image Quality
  • VIF (Sheikh & Bovik, 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
@@
  • MAD (Larson & Chandler, 2010) (Mean Absolute Deviation) provides a straightforward assessment of pixel-level differences, with lower values indicating greater similarity between images.
@@
  • CLIPIQA (Radford et al., 2021b) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
  • TReS (Radford et al., 2021b) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
@@
  • QAlign (Wu et al., 2024b) evaluates image quality by assessing its alignment with semantic and perceptual features indicative of high visual fidelity.
@@ C.3 Alignment
C.4 Aesthetic
  • NIMA (Esfandarani & Milanfar, 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images. It generates a score that reflects human perception of visual appeal.
  • LAION-AES (Schuhmann et al., 2022a) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.
  • QAlign (Wu et al., 2024b) assesses image aesthetics by analyzing its adherence to human visual preferences and artistic principles, leveraging semantic coherence and appeal.
@@ C.4 Aesthetic
Table 7: The detail of evaluation metrics.
@@
Naive Quality Pixel-level SSIM TRUE
@@
Figure 12: We evaluate nine image-generating models from 14 aspects. The numerical values in the radar chart represent the mean difference rate of each model and each metric.
Appendix D Evaluation Results
D.1 Naive Quality
@@
D.2 Applicability
We put the applicability results of Image Recognition and Semantic Segmentation in Tables 8 and  9.
Table 8: Image Recognition Discrepancy Rate ($DR_{re}$).
GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 68.60 85.46 66.14 65.50 68.12 64.88 64.88 65.36 69.26
I2I - - 94.29 94.07 40.90 83.30 27.96 47.98 -
TI2I - - 61.62 59.71 60.05 55.54 27.46 - -
Table 9: IoU of Object Segmentation.
GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 0.23 0.24 0.26 0.25 0.26 0.25 0.26 0.26 0.28
I2I - - 0.38 0.38 0.69 0.50 0.82 0.42 -
TI2I - - 0.43 0.43 0.29 0.53 0.82 - -
D.3 Human Perceptual Evaluation
D.3.1 Interface for Human Evaluation
The human evaluation of our D-ANI dataset was conducted using a custom-designed interface (Figure 14). This interface allowed participants to evaluate AI-generated images in comparison to reference content across three key aspects: Quality, Alignment, and Aesthetics. The reference content included prompts used to generate the images, categorized into three types: 1) text, 2) image, and 3) text and image. The AI-generated images evaluated were produced by four different models: SD21, SDxl, Versatile Diffusion, and DALL·E 2.
D.3.2 Evaluation Process
A total of 58 volunteers participated in the evaluation process. Each volunteer assessed the images based on the provided reference content. To ensure the validity of the data collected, we included duplicate images in the evaluation system. By comparing the scores given to these identical images, we could determine the consistency of each participant’s evaluations. Only data from participants who showed consistent scoring were considered valid, resulting in 47 effective samples.
D.3.3 Participant Demographics
The volunteers who participated in the evaluation ranged in age from 19 to 46 years old, with a gender ratio of 6:4 (male to female). All participants held at least an undergraduate degree, ensuring a well-educated sample group. Alongside scoring data, we collected demographic details such as age and gender. Upon completing their evaluations, participants were also invited to provide feedback on the evaluation system and their experience with assessing the differences between natural and AI-generated images.
D.3.4 Evaluation Metrics
Participants were instructed to evaluate each image based on three aspects:
  1. 1.
    Quality: Overall visual fidelity and clarity of the image.
  2. 2.
    Alignment: Accuracy in matching the image to the provided prompt.
  3. 3.
    Aesthetic: Visual appeal and artistic quality of the image.
To facilitate the scoring process, we provided reference examples within the evaluation system (Figure 15). These examples illustrated what constituted good or poor performance in each aspect without giving specific scores, thereby preventing any potential bias in participant evaluations.
Figure 13: Examples for unsafe references.
@@ D.3.4 Evaluation Metrics
Figure 14: Interface for Human Evaluation. Reference content: the prompt guidance used to generate images, including three types: 1) text, 2) image, and 3) text and image. AIGC images: AI-generated images that come from four different models. Participants can rely on the reference content to evaluate each image on three aspects, i.e., Quality, Alignment, and Aesthetic.
@@
Figure 15: Before the evaluation process, the participant will be given some examples to learn which image should be good or bad in a specific aspect.
@@
Figure 16: The differential rate of data from different categories models across naive image quality, alignment, and aesthetic.
@@
Figure 17: The FID, CLIP-FID and Inception Score (IS) of AI-generated images from different source models and categories from content distribution level.
D.3.5 Summary
The comprehensive human evaluation process ensured that the collected data was reliable and reflected genuine human perception. The demographic diversity and the rigorous validation of participant responses provided a robust foundation for assessing the performance of various AIGC models in generating high-quality, aligned, and aesthetically pleasing images. This approach not only highlighted the strengths and weaknesses of the evaluated models but also offered valuable insights into the perceptual differences between AI-generated and natural images.
D.4 Discrepancies Across Categories
@@
This section provides a detailed evaluation of AI-generated images across different categories, focusing on naive image quality, alignment, aesthetics, pixel-level similarity, and content distribution. These analyses aim to uncover discrepancies between AI-generated and natural images and identify category-specific challenges.
@@
Figure 17 presents trends in SSIM, FID, CLIP-FID, and Inception Score. The ‘accessory’ category consistently shows higher FID values, reflecting poorer content distribution, while the ‘animal’ category achieves higher SSIM scores, demonstrating better pixel-level similarity. These findings suggest that ‘accessory’ and ‘indoor’ images pose greater challenges for generative models in maintaining semantic consistency and achieving balanced content generation compared to other categories.
Additionally, the results in Figures 17 and Figures 16 show that compared to other generative models, images generated by D·E 3 display greater differences from natural reference images across various metrics. This might be due to D·E 3’s tendency to produce high-quality images with a surreal style, whereas other models primarily focus on generating images with a more realistic style (see Appendix Sec. B, Figure 9 for visual comparisons).
@@ Appendix E Discussion
Table 10: Prompts trigger the safe-checker of different models.
T2I I2I TI2I T2I(‰) I2I(‰) TI2I(‰)

### v3

@@ Appendix A Appendix Overview
Appendix B Additional Details of the D-ANI Dataset
The D-ANI dataset is a carefully curated collection of both natural and AI-generated images, designed to support comprehensive evaluations of image generation models. The natural images are sourced from the COCO validation set. AI-generated images were produced using nine different models. These images were generated under three distinct guidance modes: Text-to-Image (T2I), Image-to-Image (I2I), and Text vs. Image-to-Image (TI2I). Representative samples from the D-ANI dataset are shown in Figures 10,  11 and  12.
Figure 10. Examples of AI-generated content (AIGC) created using Text-to-Image (T2I) guidance.
Figure 11. Examples of AI-generated content (AIGC) created using Image-to-Image (I2I) guidance.
Figure 12. Examples of AI-generated content (AIGC) created using Text-and-Image-to-Image (TI2I) guidance.
@@ Appendix B Additional Details of the D-ANI Dataset
  • Versatile Diffusion: (Xu et al., 2023) is a diffusion-based model designed for a variety of image generation tasks, including style transfer, image inpainting, and super-resolution. Its versatility lies in its ability to adapt the diffusion process to the specific requirements of each task, making it suitable for diverse applications.
@@ C.1. Overview
To evaluate image quality, we divided the assessment into three sub-aspects: structure-level similarity, frame-level visual metrics, and visual content distribution.
  • Structure-level: Metrics such as SSIM, PSNR, LPIPS, DISTS, VIF, VSI, FSIM, and MAD assess structural, perceptual, and textural fidelity. These metrics measure how closely the generated images match natural ones in terms of structural integrity and perceptual similarity.
@@
For alignment, we utilize the CLIP Score (Hessel et al., 2021) to measure the correspondence between text descriptions and the generated images. This metric evaluates how well the generated content aligns with the given prompts by embedding both text and image into a shared space and calculating their similarity.
@@
C.2. Naive Image Quality
We evaluated the naive image quality of AI-generated images versus natural images using a range of metrics. These metrics assess both structure-level and frame-level qualities, as well as visual content distribution, to capture the holistic differences between AI-generated and natural images.
Structure-Level Quality Metrics:
@@ C.2. Naive Image Quality
  • VIF (Sheikh and Bovik, 2006) (Visual Information Fidelity) quantifies the amount of visual information preserved in the test image relative to the reference, based on natural scene statistics and the human visual system.
@@
  • MAD (Larson and Chandler, 2010) (Mean Absolute Deviation) provides a straightforward assessment of structure-level differences, with lower values indicating greater similarity between images.
@@
  • CLIPIQA (Radford et al., 2021) (CLIP Image Quality Assessment) uses CLIP model embeddings to assess image quality by comparing the alignment between image and text descriptions.
  • TReS (Radford et al., 2021) (Textural and Edge-based Similarity) evaluates image quality by focusing on textural and edge-based features, providing a no-reference assessment.
@@
  • QAlign (Wu et al., 2024) evaluates image quality by assessing its alignment with semantic and perceptual features indicative of high visual fidelity.
@@ C.3. Alignment
C.4. Aesthetic
  • NIMA (Esfandarani and Milanfar, 2018) (Neural Image Assessment) is a deep learning-based model that predicts the aesthetic quality of images. It generates a score that reflects human perception of visual appeal.
  • LAION-AES (Schuhmann et al., 2022) (LAION Aesthetic Score) evaluates images’ aesthetic quality using various visual features, providing a numerical value to indicate their aesthetic appeal.
  • QAlign (Wu et al., 2024) assesses image aesthetics by analyzing its adherence to human visual preferences and artistic principles, leveraging semantic coherence and appeal.
@@ C.4. Aesthetic
Table 9. The detail of evaluation metrics.
@@
Naive Quality Structure-level SSIM TRUE
@@
Figure 13. We evaluate nine image-generating models from 14 aspects. The numerical values in the radar chart represent the mean difference rate of each model and each metric.
Appendix D Evaluation
D.1. Naive Quality
@@
D.1.1. Aesthetic Appeal
We present the additional results of QAlign (Wu et al., 2024) regarding to the Aesthetic Appeal assessment in Table 10.
Table 10. Aesthetic Appeal of QAlign.
Types DFGAN GALIP SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 45.33 39.35 32.44 31.84 35.31 40.91 43.73 26.53 99.34
I2I - - 23.47 23.09 11.31 20.32 8.17 13.88 -
TI2I - - 25.8 25.98 23.76 26.62 7.46 - -
D.2. Applicability Settings
To facilitate the applicability evaluation in downstream tasks, we construct three standardized Visual Answering Question (VAQ) templates using BLIP-2, each targeting a specific aspect of visual-semantic understanding: object recognition, semantic alignment, and similarity probability. As shown in Table 11, each prompt is designed to elicit structured and interpretable responses. The object recognition prompt asks the model to identify the main object in the image and return a class label from the COCO taxonomy. The alignment prompt evaluates whether the image content semantically matches a given caption, requiring a binary “yes” or “no” answer. The similarity probability prompt estimates the degree of alignment between image and caption by requesting a numerical probability. These prompts are uniformly applied to both AI-generated and reference images, offering a consistent protocol for probing semantic fidelity across different generation conditions.
Table 11. The prompt template used for VAQ tasks.
Task Prompts
Object Question: what is the object in the image? Only return the class name in {COCO class list}. Answer:
Alignment Question: the image content aligns very well with {caption}? only return ’yes’ or ’no’. Answer:
Similarity probability Question: what’s the probability of the image content well align with {caption}? only return a float number. Answer:
D.3. Human Perceptual Evaluation
To complement the description in Sec.4.6, we present two figures that detail the setup of our human evaluation protocol. Figure15 illustrates the custom-designed interface through which participants evaluated AI-generated images. The interface supports side-by-side comparisons and collects ratings across three key dimensions: Quality, Alignment, and Aesthetics. For each dimension, participants assigned a score ranging from 1 to 5 using a radio-button selection mechanism, where higher scores indicated better performance. Figure 16 shows the set of reference examples embedded within the interface. These examples were designed to convey representative cases of both high and low performance for each evaluation dimension. They served as non-numeric visual guidelines to help participants develop a consistent evaluation criterion, while avoiding explicit scoring anchors that might introduce bias.
Figure 14. Examples for unsafe references.
@@ D.3. Human Perceptual Evaluation
Figure 15. Interface for Human Evaluation. Reference content: the prompt guidance used to generate images, including three types: 1) text, 2) image, and 3) text and image. AIGC images: AI-generated images that come from four different models. Participants can rely on the reference content to evaluate each image on three aspects, i.e., Quality, Alignment, and Aesthetic.
@@
Figure 16. Before the evaluation process, the participant will be given some examples to learn which image should be good or bad in a specific aspect.
@@
Figure 17. The differential rate of data from different categories models across naive image quality, alignment, and aesthetic.
@@
Figure 18. The FID, CLIP-FID, and Inception Score (IS) of AI-generated images from different source models and categories from the content distribution level.
D.4. Discrepancies Across Categories
@@
This section provides a detailed evaluation of AI-generated images across different categories, focusing on naive image quality, alignment, aesthetics, structure-level similarity, and content distribution. These analyses aim to uncover discrepancies between AI-generated and natural images and identify category-specific challenges.
@@
Figure 18 presents trends in SSIM, FID, CLIP-FID, and Inception Score. The ‘accessory’ category consistently shows higher FID values, reflecting poorer content distribution, while the ‘animal’ category achieves higher SSIM scores, demonstrating better structure-level similarity. These findings suggest that ‘accessory’ and ‘indoor’ images pose greater challenges for generative models in maintaining semantic consistency and achieving balanced content generation compared to other categories.
@@ Appendix E Discussion
Table 12. Prompts trigger the safe-checker of different models.
Models T2I I2I TI2I T2I(‰) I2I(‰) TI2I(‰)